# sDPO: Don’t Use Your Data All at Once

### v1

@@ Abstract
As development of large language models (LLM) progresses, aligning them with human preferences has become increasingly important. We propose stepwise DPO (sDPO), an extension of the recently popularized direct preference optimization (DPO) for alignment tuning. This approach involves dividing the available preference datasets and utilizing them in a stepwise manner, rather than employing it all at once. We demonstrate that this method facilitates the use of more precisely aligned reference models within the DPO training framework. Furthermore, sDPO trains the final model to be more performant, even outperforming other popular LLMs with more parameters.
@@
Dahyun Kim, Yungi Kim, Wonho Song, Hyeonwoo Kim, Yunsu Kim, Sanghoon Kim Chanjun Park Upstage AI, South Korea {kdahyun, eddie, ynot, choco_9966, yoonsoo, limerobot, chanjun.park}@upstage.ai
@@ 1 Introduction
Large language models (LLMs) have revolutionized the field of natural language processing (NLP) through a training process that includes pre-training, supervised fine-tuning, and alignment tuning, with the latter ensuring the safety and usefulness of the model. Thus, reinforcement learning techniques Christiano et al. (2017); Bai et al. (2022), such as proximal policy optimization (PPO) Schulman et al. (2017), are key in this alignment phase, despite their complexity.
To address the complicated nature of reinforcement learning in LLM training, direct preference optimization (DPO) Rafailov et al. (2023), among other methods Yuan et al. (2023); Dong et al. (2023), have been popularized for its simplicity and effectiveness. DPO involves curating preference datasets using human or strong AI (e.g., GPT-4 OpenAI (2023)) judgement to select chosen and rejected responses to questions. These datasets are used to train LLMs by comparing log probabilities of chosen versus rejected answers. However, obtaining these probabilities can be challenging with proprietary models like GPT-4, since they do not offer log probabilities for inputs.
@@
Thus, in most practical scenarios, the reference model is simply set as the base SFT model Tunstall et al. (2023); Intel (2023b); Ivison et al. (2023), which is a much weaker alternative with potentially misaligned preferences. This reference model acts as a lower bound in DPO, i.e., the target model is optimized to be at least as aligned as the reference model. Thus, we argue that a reference model that is already more aligned will serve as a better lower bound for DPO training, which would be beneficial for the alignment tuning. One option would be to utilize the plethora of open source models Tunstall et al. (2023); Ivison et al. (2023) that have already undergone alignment tuning.
@@
Note that the above may not be feasible due to the absence of such aligned models, or the fact that it renounces control over the reference model, leading to safety concerns. Instead, we propose ‘stepwise DPO’, named sDPO, where we use the preference datasets (or subsets of a preference dataset) in a step-by-step manner when undergoing DPO training. The aligned model in the previous step is used as the reference model for the current step, which results in utilizing a more aligned reference model (i.e., a better lower bound). Empirically, we show that using sDPO results in a more performant final aligned model as well.
@@ 2.1 Preliminary Investigation on Reference Models
To gauge the importance of using a well-aligned reference model in DPO, we perform preliminary experiments of DPO training with the Ultrafeedback dataset Cui et al. (2023) on Mistral-7B-OpenOrca Lian et al. (2023) and OpenHermes-2.5-Mistral-7B Teknium (2023) as the SFT base model, owing to their excellent performance and small size. We compare the following reference models: i) the SFT base model itself, same as the conventional DPO setup; ii) SOLAR-0-70B Upstage (2023), a larger and much more performant model; and iii) Intel-7B-DPO Intel (2023a), an already aligned reference model. The results are summarized in Tab. 1.
As the table shows, using Intel-7B-DPO as the reference model results in the best performance, even better than using SOLAR-0-70B, which is a much larger model that was trained with more data. Thus, whether the reference model is pre-aligned or not plays an important role in the resulting aligned model’s performance. Unfortunately, it is not always possible to simply use a open sourced pre-aligned model as the reference model due to technical and safety concerns, i.e., such a model may not exist yet or can be susceptible to various domain-specific harmfulness and fairness criteria. To remedy the above, we propose sDPO, which uses more aligned reference models as a part of the training framework.
@@ 2.2 Stepwise DPO
In sDPO, we propose to use the available preference datasets in a stepwise manner instead of using them all at once. The comparison of the overall flow of DPO and sDPO is presented in Fig. 1.
@@
Table 2: Performance comparison of applying sDPO (and ablated versions) to SOLAR 10.7B + SFT against various top performing models. Size is shown in units of billions of parameters and type is reported as one of {‘Pretrained’, ‘Instruction-tuned’, ‘Alignment-tuned’}. Models based on SOLAR 10.7B are shown in purple color. The best scores in each column are shown in bold.
Intuitive explanation.
@@
where $D$ is the preference dataset, $x$ is the question, $y_{w}$ and $y_{l}$ are the chosen and rejected answers respectively, $\theta$ is the learnable parameters of the model, and $\gamma_{\pi}(x,y_{w},y_{l})=\log{{\pi(y_{w}|x)\over\pi(y_{l}|x)}}$, i.e., the logratio of the chosen and rejected samples w.r.t. the policy $\pi$. As $\log\sigma(\cdot)$ is a monotonically increasing function and $\gamma_{\pi_{ref}}$ is fixed before training, the minimization of $\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{ref})$ leads to $\gamma_{\pi_{\theta}}>\gamma_{\pi_{ref}}$ (on average). Thus, $\gamma_{\pi_{ref}}$ can be understood as a lower bound defined by the reference model, of which the target model is trained such that $\gamma_{\pi_{\theta}}>\gamma_{\pi_{ref}}$. In sDPO, $\gamma_{\pi_{ref}}$ increases as the steps progress because the reference model that defines it is more and more aligned. Hence, $\gamma_{\pi_{ref}}$ becomes a stricter lower bound as the steps pass, inducing a curriculum learning from easy to hard optimization tasks.
@@ 3.1 Experimental Setup
We use a supervised fine-tuned SOLAR 10.7B Kim et al. (2023) as our SFT base model $S$ as it delivers excellent performance with its uncommon 10.7B size. Further, the scarcity of 10.7B sized models leads to the absence of open source models that can be adopted as reference models, making the usage of sDPO more necessary. We use OpenOrca Mukherjee et al. (2023) ($\sim 12K$ samples) and Ultrafeedback Cleaned ($\sim 60K$ samples) Cui et al. (2023); Ivison et al. (2023) as our preference datasets. The training hyper-parameters closely follow that of Tunstall et al. (2023). We use two steps in sDPO, where we use OpenOrca as dataset $D_{1}$ in the first step and Ultrafeedback Cleaned as dataset $D_{2}$ in the second step.
@@
We utilize four of the six tasks in the HuggingFace Open LLM Leaderboard Beeching et al. (2023): ARC Clark et al. (2018), HellaSWAG Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2022). We also report the average scores for the four tasks, which is denoted as H4. Winogrande Sakaguchi et al. (2021) and GSM8K Cobbe et al. (2021) are excluded to control the complexity of the experiments, i.e., we excluded generation tasks in contrast to multiple choice tasks.
@@ 3.2 Main Results
Evaluation results for applying sDPO to the SFT base model, along with results for other top-performing models are shown in Tab. 2. Comparing the pretrained-only ‘SOLAR 10.7B’ to the instruction-tuned ‘SOLAR 10.7B + SFT’, we can see an increase of $+5.24$ in terms of H4. Applying sDPO on SOLAR 10.7B + SFT further increases the H4 score upto $74.31$, an improvement of $+4.80$. Notably, ‘SOLAR 10.7B + SFT + sDPO’ outperforms other larger models such as Mixtral 8x7B-Instruct-v0.1, despite the smaller number of parameters. This highlights that effective alignment tuning could be the key to unlocking next level performance for smaller LLMs. Further, applying sDPO results in substantially higher score of $72.45$ for TruthfulQA, which shows the effectiveness of the alignment tuning process.
3.3 Ablation Studies
We also report evaluation results for ablated models in Tab. 2. ‘SOLAR 10.7B + SFT + DPO’ uses all the DPO data at once, i.e., $D_{1}+D_{2}$, same as the conventional DPO training setup. ‘SOLAR 10.7B + SFT + sDPO Strat.’ uses stratified sampling to sample $\sim 16.67\%$ of the data points from the union of OpenOrca and Ultrafeedback Cleaned to form $D_{1}$ and use the remaining $\sim 83.33\%$ as $D_{2}$ to mirror the dataset size of $D_{1}$ and $D_{2}$ used in SOLAR 10.7B + SFT + sDPO.
Comparing SOLAR 10.7B + SFT + DPO and SOLAR 10.7B + SFT + sDPO, we can see that using sDPO over DPO results in a higher H4 score overall, with noticeable improvements in ARC and TruthfulQA scores. Therefore, we believe sDPO could function as a drop-in replacement for DPO training with better performance. Looking at SOLAR 10.7B + SFT + sDPO and SOLAR 10.7B + SFT + sDPO Strat., we see that the specific way of splitting the available DPO data into multiple $D_{t}$ can also impact performance. We find that the natural split of using different preference datasets as $D_{t}$ works best in our experiments. We believe further exploration of how to define $D_{t}$ is an interesting direction for future research.
@@ 3.4 Reference Models in sDPO
In Sec. 2.2, we explain that the reference models in sDPO are more aligned, resulting in higher $\gamma_{\pi_{ref}}$, i.e., a stricter lower bound. We verify the above empirically in Fig. 2 by comparing the mean $\gamma_{\pi_{ref}}$ on the Ultrafeedback Cleaned dataset for the reference models in steps 1 and 2 of sDPO, i.e., $S$ and $M_{1}$. Note that these two models have not been trained on the aforementioned dataset. Using the SFT base model $S$ as the reference model, the mean of $\gamma_{\pi_{ref}}$ is $-38.60$. On the other hand, using the aligned model $M_{1}$ from step 1 of sDPO as the reference model, the mean of $\gamma_{\pi_{ref}}$ is $-25.10$, an increase of $13.50$ in log scale. Thus, a single step of sDPO greatly increases $\gamma_{\pi_{ref}}$, which results in a more performant aligned model as seen in Tab. 2.
@@
We also show mean $\gamma_{\pi_{ref}}$ of $M_{2}$, the aligned model from step 2 of sDPO. Unlike $S$ and $M_{1}$, $M_{2}$ is trained on the Ultrafeedback Cleaned dataset, i.e., $M_{2}$ is used as a reference model on data that was already used to train it. Note that such a case could happen commonly when adopting various open source models as reference models. This is because the datasets that were used in training those models are often unclear and could overlap with the preference datasets unintentionally. Mean $\gamma_{\pi_{ref}}$ of $M_{2}$ is $84.35$, which is staggeringly higher than either $S$ or $M_{1}$. The strikingly high value for $M_{2}$ likely points to overfitting of $M_{2}$ to the Ultrafeedback Cleaned dataset. This result highlights the potential danger of merely adopting open source models as reference models instead of using sDPO.

### v2

@@ Abstract
As large language models (LLMs) continue to advance, aligning them with human preferences has become a critical objective. In this paper, we introduce stepwise DPO (sDPO), an innovative extension of the recently popularized Direct Preference Optimization (DPO) technique for alignment tuning. sDPO systematically partitions the available preference datasets and applies them incrementally, rather than utilizing the entire dataset simultaneously. This stepwise manner enables the integration of progressively more aligned reference models within the DPO training framework. Our empirical results demonstrate that sDPO not only enhances the alignment precision of reference models but also significantly improves the overall performance of the final model, surpassing other prominent LLMs with larger parameter counts.
@@
Dahyun Kim, Yungi Kim, Wonho Song, Hyeonwoo Kim, Yunsu Kim, Sanghoon Kim Chanjun Park Upstage AI {kdahyun, eddie, ynot, choco_9966, yoonsoo, limerobot, chanjun.park}@upstage.ai
@@ 1 Introduction
Large language models (LLMs) have revolutionized the field of natural language processing (NLP) by undergoing pre-training, supervised fine-tuning, and alignment tuning, with the latter ensuring the safety and usefulness of the model. Reinforcement learning (RL) techniques Christiano et al. (2017); Bai et al. (2022), such as proximal policy optimization (PPO) Schulman et al. (2017), are generally used in this alignment phase.
To address the complicated nature of RL in LLM training, direct preference optimization (DPO) Rafailov et al. (2023) have been popularized for its simplicity and effectiveness. DPO involves curating preference datasets using human or strong AI (e.g., GPT-4 OpenAI (2023)) judgement to select chosen and rejected responses from a pool of multiple answers to a given question. Then, the model being trained (i.e., target model) and a separate reference model compute log probabilities of chosen and rejected responses. Finally, the target model is trained by maximizing the difference of the log probability ratios of the target and the reference models for the chosen and rejected answers. However, obtaining these probabilities can be challenging if one wants to use proprietary models like GPT-4 as the reference model, since they do not offer log probabilities for inputs.
@@
Thus, in practice, the reference model is simply set as the base SFT model Tunstall et al. (2023); Intel (2023b); Ivison et al. (2023), which is a much weaker alternative with potentially misaligned preferences. Through Eq. 1, we show that the reference model acts as a lower bound in DPO, i.e., the target model is optimized to be at least as aligned as the reference model. Thus, we argue that a reference model that is already more aligned will serve as a better lower bound for DPO training, which would be beneficial for the alignment tuning. One option would be to utilize the plethora of open source models Tunstall et al. (2023); Ivison et al. (2023) that have already undergone alignment tuning.
@@
Note that the above approach may not be feasible due to the absence of such aligned models, or the fact that it renounces control over the reference model, which could lead to safety concerns. Instead, we propose ‘stepwise DPO’, named sDPO, where we use the preference datasets (or subsets of a preference dataset) in a step-by-step manner rather than all at once when undergoing DPO training. The aligned model in the previous step is used as the reference model for the current step, which results in utilizing a more aligned reference model (i.e., a better lower bound). Empirically, we show that using sDPO results in a more performant final aligned model as well.
@@
2 Related Work
2.1 Large Language Models
Recent research has highlighted a "scaling law" in the field of context-based language models Kaplan et al. (2020); Hernandez et al. (2021); Anil et al. (2023), showing a proportional relationship between the size of the model plus the training data and the resulting performance improvements. Consequently, this has led to the advent of LLMs. In contrast to earlier models, LLMs can perform in-context learning, which includes abilities such as zero-shot learning Radford et al. (2019) and few-shot learning Brown et al. (2020), allowing them to adapt and perform tasks without the need for weight adjustments. These emergent abilities of LLMs, absent in their smaller counterparts, signal a significant evolution in language model capabilities Wei et al. (2022).
2.2 Alignment Tuning
To address these issues, there have been proposals for supervised fine-tuning methodologies such as RRHF Yuan et al. (2023), RAFT Dong et al. (2023), and DPO Rafailov et al. (2023). These methods circumvent the intricacies inherent in reinforcement learning and have been shown to yield empirical results on par with RLHF. Notably, the DPO technique straightforwardly encourages the LLM to favor positive responses and discourage negative ones. DPO has been observed to yield performant learning outcomes, in spite of its uncomplicated training procedure.
Concurrent to our work, Yuan et al. (2024) have developed an iterative framework for generating new preference datasets and performing DPO training on the resulting datasets. They empirically demonstrated the superiority of their iterative framework in terms of AlpacaEval 2.0. In contrast, our work is complementary to the above in the sense that we focus on utilizing the current preference data and does not undergo new data generation. Thus, our method can also be applied to Yuan et al. (2024) by changing the DPO training part to using sDPO instead. Additionally, the evaluation used in Yuan et al. (2024) is also different to ours as we utilize tasks from Open LLM Leaderboard Beeching et al. (2023), EQ Bench Paech (2023) and MT Bench Zheng et al. (2023) whereas Yuan et al. (2024) uses AlpacaEval 2.0.
@@ 3.1 Preliminary Investigation on Reference Models
To gauge the importance of using a well-aligned reference model in DPO, we perform preliminary experiments of DPO training with the Ultrafeedback dataset Cui et al. (2023) on Mistral-7B-OpenOrca Lian et al. (2023) and OpenHermes-2.5-Mistral-7B Teknium (2023) as the SFT base model, owing to their excellent performance and small size. We compare the following reference models: i) the SFT base model itself, same as the conventional DPO setup; ii) SOLAR-0-70B Upstage (2023), a larger and much more performant model; and iii) Intel-7B-DPO Intel (2023a), an already aligned reference model. The results are summarized in Table 1.
As the table shows, using Intel-7B-DPO as the reference model results in the best performance, even better than using SOLAR-0-70B, which is a much larger and performant model. Thus, whether the reference model is pre-aligned or not plays an important role in the resulting aligned model’s performance. Unfortunately, it is not always possible to use a open sourced pre-aligned model as the reference model due to technical and safety concerns. For instance, such a model may not exist yet or can be susceptible to various domain-specific harmfulness and fairness criteria along with potential data contamination issues. To circumvent the above, we propose sDPO, which does not require an external pre-aligned model but uses more aligned reference models, built from the SFT base model, as a part of the training framework.
@@ 3.2 Stepwise DPO
In sDPO, we propose to use the available preference datasets in a stepwise manner instead of using them all at once. Essentially, we partition the preference data into $T$ chunks and perform DPO training $T$ times. The trained model from the previous step is used as the reference and target models, which means that each of the $T$ DPO training steps function in a similar manner to the conventional DPO setup. In doing so, we create and utilize intermediary reference models that are more aligned than those thar are used in conventional DPO. The comparison of the overall flow of DPO and sDPO is presented in Figure 1.
@@
Table 2: Performance comparison of applying sDPO or DPO to SOLAR 10.7B + SFT against various top performing models. Size is shown in units of billions of parameters and type is reported as one of {‘Pretrained’, ‘Instruction-tuned’, ‘Alignment-tuned’}. Models based on SOLAR 10.7B are shown in purple color. The best scores in each column are shown in bold.
Mathematical explanation.
@@
where $D$ is the preference dataset, $x$ is the question, $y_{w}$ and $y_{l}$ are the chosen and rejected answers respectively, $\theta$ is the learnable parameters of the model, and $\gamma_{\pi}(x,y_{w},y_{l})=\log{{\pi(y_{w}|x)\over\pi(y_{l}|x)}}$, i.e., the logratio of the chosen and rejected samples w.r.t. the policy $\pi$. As $\log\sigma(\cdot)$ is a monotonically increasing function and $\gamma_{\pi_{ref}}$ is fixed before training, the minimization of $\mathcal{L}_{\text{DPO}}(\pi_{\theta},\pi_{ref})$ leads to $\gamma_{\pi_{\theta}}>\gamma_{\pi_{ref}}$ on average. Thus, $\gamma_{\pi_{ref}}$ can be understood as a lower bound defined by the reference model, of which the target model is trained such that $\gamma_{\pi_{\theta}}>\gamma_{\pi_{ref}}$. In sDPO, $\gamma_{\pi_{ref}}$ increases as the steps progress because the reference model that defines it is more and more aligned. Hence, $\gamma_{\pi_{ref}}$ becomes a stricter lower bound as the steps pass, inducing a curriculum learning from easy to hard optimization tasks. Thus, the target model is being trained to learn stricter preferences as the steps progress in sDPO.
Data partitioning strategy.
The method for partitioning the preference data into $T$ chunks is also important in sDPO. One option would be to pool all the data from different dataset sources and perform random sampling. However, we argue that partitioning the data such that earlier chunks are comprised of easier preference data would be more aligned with inducing a curriculum learning of easy to hard optimization in sDPO.
To that end, we propose to use easy-to-hard data partitioning by the following method. Using $M_{0}$, the initial target model, we calculate the reward accuracy, i.e., the percentage of samples in which the target model scores higher rewards for preferred samples, for the different dataset sources. The dataset sources are sorted in descending order of the reward accuracy, which are then used as the $T$ chunks in sDPO. Thus, if we have $N$ dataset sources, we would have a total of $N$ chunks, where earlier chunks would contain easier samples as measured by the reward accuracy.
@@ 4.1 Experimental Setup
We use a supervised fine-tuned SOLAR 10.7B Kim et al. (2023) as our SFT base model $S$ as it delivers excellent performance with its uncommon yet relatively small 10.7B size. Note that we do not need a separate reference model as it is initialized as $M_{t-1}$, the final trained model from the previous step. We use OpenOrca Mukherjee et al. (2023) ($\sim 12K$ samples) and Ultrafeedback Cleaned ($\sim 60K$ samples) Cui et al. (2023); Ivison et al. (2023) as our preference datasets. The training hyper-parameters follow that of Tunstall et al. (2023). Using the easy-to-hard partitioning, we use OpenOrca as dataset $D_{1}$ and Ultrafeedback Cleaned as dataset $D_{2}$.
@@
We mainly utilize four log-probability tasks in the HuggingFace Open LLM Leaderboard Beeching et al. (2023): ARC Clark et al. (2018), HellaSWAG Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2022). We also report the average scores for the four tasks, which is denoted as H4. Note that these tasks do not require the model to actually generate a new answer to the question. Rather, the log-probability of a pre-defined answer is measured instead.
To augment the above potential downside of log-probability benchmarks, we also incorporate generation benchmarks such as EQ Bench Paech (2023) and MT Bench Zheng et al. (2023), where a model is prompted to generate an answer to a question. As such, MT Bench and EQ Bench both strongly correlate with the Chatbot Arena ELO Zheng et al. (2023); Chiang et al. (2024), one of the most widely recognized open-world LLM evaluation system.
@@ 4.2 Main Results
Evaluation results for applying sDPO to the SFT base model, along with results for other top-performing models are shown in Table 2. Applying sDPO on SOLAR 10.7B + SFT further increases the H4 score to $74.31$, an improvement of $+4.80$. Notably, ‘SOLAR 10.7B + SFT + sDPO’ outperforms other larger models such as Mixtral 8x7B-Instruct-v0.1, despite the smaller number of parameters. This highlights that effective alignment tuning could be the key to unlocking next level performance for smaller LLMs. Further, applying sDPO results in substantially higher score of $72.45$ for TruthfulQA, which demonstrates the effectiveness of the alignment tuning process. We also present additional results in Table 4 of Section 4.7 on the EQ Bench Paech (2023), which is a generation task with high correlation with the Chatbot Arena ELO Zheng et al. (2023). The additional results indicate the superiority of sDPO over DPO in improving generation task performance as well.
4.3 Ablation Studies Against DPO
We also report evaluation results for ablating sDPO with traditional DPO in Table 2. ‘SOLAR 10.7B + SFT + DPO’ uses all the DPO data at once, i.e., $D_{1}+D_{2}$, same as the conventional DPO training setup.
We can see that using sDPO over DPO results in a higher H4 score overall, with noticeable improvements in ARC and TruthfulQA scores. Therefore, we believe sDPO could function as a drop-in replacement for DPO training with better performance.
@@ 4.4 Reference Models in sDPO
In Sec. 3.2, we explain that the reference models in sDPO are more aligned, resulting in higher $\gamma_{\pi_{ref}}$, i.e., a stricter lower bound. We verify the above empirically in Figure 2 by comparing the mean $\gamma_{\pi_{ref}}$ on the Ultrafeedback Cleaned dataset for the reference models in steps 1 and 2 of sDPO, i.e., $S$ and $M_{1}$. Note that these two models have not been trained on the aforementioned dataset. Using the SFT base model $S$ as the reference model, the mean of $\gamma_{\pi_{ref}}$ is $-38.60$. On the other hand, using the aligned model $M_{1}$ from step 1 of sDPO as the reference model, the mean of $\gamma_{\pi_{ref}}$ is $-25.10$, an increase of $13.50$ in log scale. Thus, a single step of sDPO greatly increases $\gamma_{\pi_{ref}}$, which results in a more performant aligned model as seen in Table 2.
@@
We also show mean $\gamma_{\pi_{ref}}$ of $M_{2}$, the aligned model from step 2 of sDPO. Unlike $S$ and $M_{1}$, $M_{2}$ is trained on the Ultrafeedback Cleaned dataset, i.e., $M_{2}$ is used as a reference model on data that was already used to train it. Note that such a case could happen commonly when adopting various open source models as reference models. This is because the datasets that were used in training those models are often unclear and could overlap with the preference datasets unintentionally. Mean $\gamma_{\pi_{ref}}$ of $M_{2}$ is $84.35$, which is staggeringly higher than either $S$ or $M_{1}$. The strikingly high value for $M_{2}$ likely points to overfitting of $M_{2}$ to the Ultrafeedback Cleaned dataset. Note that utilizing such an absurdly high value of $\gamma_{\pi_{ref}}$ as the lower bound in DPO training may be undesirable. This result highlights the potential danger of merely adopting open source models as reference models instead of using sDPO.
Model H4 (Avg.) ARC HellaSwag MMLU TruthfulQA
SOLAR 10.7B + SFT + sDPO 74.31 71.33 88.08 65.39 72.45
SOLAR 10.7B + SFT + sDPO Rand. 72.56 69.20 87.27 65.96 67.81
Table 3: Comparison between the easy-to-hard and random partitioning strategies. ‘SOLAR 10.7B + SFT + sDPO’ uses the easy-to-hard partitioning whereas ‘SOLAR 10.7B + SFT + sDPO Rand.’ denotes sDPO with random partitioning instead. Easy-to-hard partitioning is better than random partitioning. The best scores are shown in bold.